OpenStudy (cheesecakekitten):

@ganeshie8 @phi @lordhelix8th Which property describes this expression? a + (b + c) = (a + b) + c A. Distributive Property B. Commutative Property of Addition C. Definition of Subtraction D. Associative Property of Addition

Mehek (mehek14):

which one do you think?

OpenStudy (cheesecakekitten):

A

OpenStudy (cheesecakekitten):

Or d

Mehek (mehek14):

this is distributive: \(\bf{a(b+c)=ab+ac}\)

OpenStudy (cheesecakekitten):

I know it's not C

Mehek (mehek14):

yes it's D associative \(\bf{(a+b)+c=a+(b+c)}\)

OpenStudy (cheesecakekitten):

sorry ganeshie your too late

ganeshie8 (ganeshie8):

good to see you figured out :)